OpenStudy (haleyelizabeth2017):

Honors Algebra 2, continued. (sorry about that, computer is acting up)

OpenStudy (haleyelizabeth2017):

Divide and simplify if possible. \(\frac{ \sqrt {250x^{16}} }{ \sqrt {2x} }\)

OpenStudy (anonymous):

sorry im new to this and that equation tool isn't working for me. but you can split the numerator into sqrt(125 x^8) * sqrt(2x) * sqrt(x^7)

OpenStudy (haleyelizabeth2017):

okay....

OpenStudy (haleyelizabeth2017):

\[\frac{ \sqrt {125x^8} \sqrt {2x} \sqrt {x^7} }{ \sqrt {2x} }\]

OpenStudy (anonymous):

Yes, and then cancel the sqrt(2x) from the numerator & denominator; so you'd end up with sqrt(125x^8) * sqrt(x^7). You can multiply what's inside the square roots to get (sqrt(125x^15)) and then simplify that to = (x^7)(sqrt(125x))

OpenStudy (anonymous):

Sorry, it looks really messy typed out..

OpenStudy (haleyelizabeth2017):

okay......haha

OpenStudy (anonymous):

haha hope it was all understandable :)

OpenStudy (haleyelizabeth2017):

so \[\sqrt {125x^8} \sqrt {x^7}\] \(\sqrt {125x^{15}}\) \(x^7 \sqrt {125x}\)

OpenStudy (anonymous):

Yes :)

OpenStudy (haleyelizabeth2017):

okay thank you!

OpenStudy (anonymous):

you're welcome

OpenStudy (haleyelizabeth2017):

@EricWise here is the other method :)

OpenStudy (haleyelizabeth2017):

jacquelyn made it easy to understand :) easier at least